### v2

# Supernovae in 2023 (review): breakthroughs by late observations

Abstract
I present a review of how late observations of supernovae, of the nebular phase, and much later of supernova remnants (SNRs), and their analysis in 2023 made progress towards breakthroughs in supporting the jittering jets explosion mechanism (JJEM) for core-collapse supernovae (CCSNe) and in introducing the group of lonely white dwarf (WD) scenarios for type Ia supernovae (SNe Ia). The new analyses of CCSN remnants (CCSNRs) reveal point-symmetric morphologies in a way unnoticed before in three CCSNRs. Comparison to multipolar planetary nebulae that are shaped by jets suggests that jets exploded these CCSNe, as predicted by the JJEM, but incompatible with the prediction of the delayed neutrino explosion mechanism. The spherical morphology of the ejecta Pa 30 of the historical type Iax supernova (SN Iax) of 1181 AD, which studies in 2023 revealed, is mostly compatible with the explosion of a lonely WD. Namely, at the explosion time, there is only a WD, without any close companion, although the WD was formed via a close binary interaction, i.e., binary merger. An identification of point-symmetry in SNR G1.9+0.3, a normal SN Ia and the youngest SN in the Galaxy, suggests an SN explosion of a lonely WD inside a planetary nebula (an SNIP). The group of lonely WD scenarios includes the core degenerate scenario and the double degenerate scenario with a merger to explosion delay (MED) time. SN Ia explosions of lonely WDs are common, and might actually account for most (or even all) normal SNe Ia.
keywords:
Supernovae: general – ISM: supernova remnants – stars: massive – stars: neutron – stars: white dwarfs – stars: jets
1 Introduction
The objective of this review is to present the topics that supernova (SN) studies cover in 2023 in relation to the origin and explosion mechanisms of SNe. I review both core-collapse SNe (CCSNe; Section 3) and type Ia SNe (SNe Ia; Section 4). I emphasize some challenges that new studies in 2023 present and how these studies advance our knowledge toward possible breakthroughs.
As such, this review focuses on papers from the year 2023 that are relevant to either the explosion mechanism of CCSNe or to the correspondence between different scenarios of SNe Ia and observed types of SNe Ia, including normal SNe Ia and peculiar SNe Ia. I will not review the basic physics and observations of these supernovae, nor will I review pair-instability SNe (e.g., Schulze et al. 2023). I will almost not cite papers before the year 2023 as these are cited in the relevant papers from 2023 that I cite. For extended reviews of CCSN explosion mechanisms, the reader should consult earlier reviews, e.g., Janka et al. (2016) and Mezzacappa (2023) for reviews of the physics of CCSNe and the delayed-neutrino explosion mechanism and Soker (2022) on the jittering jets explosion mechanism (JJEM). In Soker (2022) I argued that models that attribute the extra energy of super-luminous CCSNe in the frame of the delayed neutrino explosion mechanism to magnetars must include jets because jets during and/or after the explosion accompany the formation of energetic magnetars. I will return to this important point in Section 3.1.
The different scenarios of SNe Ia motivated many reviews on the topic of SNe Ia in the last decade which together cover all aspects of SNe Ia prior to 2023 (e.g., Maoz et al. 2014; Maeda, & Terada 2016; Hoeflich 2017; Livio & Mazzali 2018; Soker 2018, 2019; Wang 2018; Jha et al. 2019; Ruiz-Lapuente 2019; Ruiter 2020; Liu, Röpke, & Han 2023). Vinkó, Szalai, & Könyves-Tóth (2023) present a short review of some recent aspects of SNe Ia and CCSNe. For observational classification see, e.g., Aleo et al. (2023).
CCSNe and SNe Ia differ in many aspects from each other. They also have some common physical processes. The common ground of CCSNe and SNe Ia in this review is the usage of supernova remnants (SNRs). I will use the morphologies of some SNRs to reach what I consider to be significant steps towards breakthroughs in 2023. One breakthrough is in coming closer to answering the question of what is the explosion mechanism of CCSNe? (Section 3.2) and the other breakthrough is in establishing the group of lonely white dwarf (WD) scenarios of SNe Ia, which I argue hints at the dominant scenarios of normal SN Ia (Section 4.2). I start (Section 2) by describing the point-symmetric morphology that is relevant to the SNRs that I describe. I open each of the following two Sections, Section 3 on CCSNe and Section 4 on SNe Ia, with a review of relevant papers from 2023 and then present new results that I claim might be considered breakthroughs. I summarize this review in Section 5.
2 The point-symmetric morphology of SNRs
Because the study of SNR morphology is relevant both to CCSNe and to the progenitors of SNe Ia, I discuss the point-symmetry morphology here.
The study of SNRs (their properties, their progenitors, their interaction with the ISM, particle acceleration, and more) is a subject by itself that continued to be a major topic in 2023 (e.g., Araya 2023; Bakış et al. 2023; Ball et al. 2023; Bamba et al. 2023; Banovetz et al. 2023; Bozzetto et al. 2023; Braun et al. 2023; Bykov et al. 2023; Caldwell & Raymond 2023; Cao et al. 2023; Chambery et al. 2023; Deng et al. 2023; Dokara et al. 2023; Domček et al. 2023; Duvidovich & Petriella 2023; Eagle et al. 2023; Ellien, Greco, & Vink 2023; Ferrazzoli et al. 2023; Fujishige et al. 2023; Ghavamian et al. 2023; Godinaud et al. 2023; Himono et al. 2023; Holland-Ashford, Slane, & Long 2023; Katsuda 2023; Khabibullin et al. 2023; Kim, Koo, & Onaka 2023; Koplitz et al. 2023; Makarenko et al. 2023; Meyer et al. 2023; Narita et al. 2023; Nguyen-Dang et al. 2023; Ranasinghe & Leahy 2023; Rho et al. 2023a, b; Sankrit, Blair, & Raymond 2023; Sano et al. 2023; Seward & Points 2023; Shanahan et al. 2023; Sharma et al. 2023a; Sinitsyna & Sinitsyna 2023; Suherli et al. 2023; Sushch & Brose 2023; Suzuki et al. 2023; Vasiliev & Shchekinov 2023; Winkler, Long, & Blair 2023; Winkler et al. 2023; Xiao et al. 2023; Xin & Guo 2023; Yamauchi & Pannuti 2023; Yeung, Bamba, & Sano 2023; Zhang et al. 2023b, c; Zhang & Xin 2023; Zhong et al. 2023; Zhou et al. 2023; Alsaberi et al. 2024; López-Caraballo et al. 2024; Reyes-Iturbide et al. 2024). I will focus only on the morphologies of some SNRs.
A point symmetric morphology is one where for each of the major morphological features in a nebula or a remnant there is an opposite morphological feature with respect to the center. Such morphological features might be clumps, filaments, lobes, and ears. An ear is a protrusion from the main nebula or remnant that has a base smaller than the main nebula/remnant and a cross-section that decreases outwards. Cotton et al. (2023) argue that ears are common in SNRs. This is also evident from the classification of CCSN remnants (CCSNRs; Soker 2023d).
The point-symmetry morphology is the main SNR property that was used in 2023 to point to a possible breakthrough in our understanding of CCSN explosion mechanisms and to establish the group of lonely WD SN Ia scenarios. I make use of the large body of knowledge that exists in the study of bipolar and point-symmetric planetary nebulae, like multi-polar planetary nebulae. Many studies of planetary nebulae show that bipolar and point-symmetric structures are shaped by jets (e.g., Akashi 2023; Gómez-Muñoz et al. 2023; Lora et al. 2023; Mari, Gonçalves, & Akras 2023; Moraga Baez et al. 2023 for some 2023 studies). Comparison of some SNR morphologies to point-symmetric planetary nebulae can lead to the identification of point-symmetric SNRs (e.g., Soker 2023d, 2024c).
The interaction of SN ejecta with a circumstellar material (CSM) or with the interstellar medium (ISM) cannot shape a point-symmetry morphology unless the ejecta and/or the CSM have point-symmetric morphology, to begin with; I here mention only studies from 2023 (see discussion in Soker 2024a). A magnetized medium cannot form a point-symmetry with three or more pairs of opposite ears/clumps/filaments/lobes as they form a structure with one symmetry axis, or rarely two symmetry axes, one of the exploding jets and one of the ambient magnetic field direction (e.g., Meyer et al. 2024). Velázquez et al. (2023) simulated a non-spherical wind interaction with a magnetized ISM. They find that in a pre-explosion axisymmetric wind with an axis inclined to the interstellar medium (ISM) magnetic field, the ears of the SNR might be bent. Namely, not exactly on a straight line through the center. However, this process cannot shape point-symmetric filaments/clumps/ears/lobes. Zhang et al. (2023d) included magnetic fields and ISM density gradients in their simulation of SNR G1.9+0.3 (that I study in Section 4.2.3). They could form an SNR with a pair of ears but not a point-symmetric SNR. Villagran et al. (2024) performed magneto-hydrodynamic simulations aiming at shaping SNR G1.9+0.3. They blew a non-spherical pre-explosion wind into a magnetized ISM. They obtained an axisymmetrical morphology, but not a point-symmetric morphology. Instabilities that are excited by ejecta-ISM interaction, as Mandal et al. (2023) study, are unable to form point-symmetric morphology. Turbulence in the interstellar medium (ISM), e.g., Rigon & Inoue (2023), also cannot form point-symmetric structures.
My conclusion is that point-symmetric SNRs are shaped by jets. In this review, I present results from 2023 attributing the point-symmetric structures of three CCSNRs (N63A, Vela, SN 1987A) to point-symmetric ejecta shaped by jets that were part of the pairs of jets that exploded the star in the frame of the JJEM, while a point-symmetric SNR Ia (SNR G1.9+0.3) is attributed to a point-symmetric CSM which was a planetary nebula.
3 Core collapse supernova ejecta
3.1 Neutrino-driven versus jet-driven explosions
I start by reviewing some (but not all) studies of CCSNe that are relevant to the quest for the explosion mechanism. In comparing theory with observations, one must bear in mind the large diversity of CCSNe, as studies in 2023 continued to show (e.g., Agudo et al. 2023; Ailawadhi et al. 2023; Aryan et al. 2023; Ben-Ami et al. 2023; Blanchard et al. 2023; Bostroem et al. 2023a, c; Chen et al. 2023b; Corsi et al. 2023; Desai et al. 2023a; Dong et al. 2023b; Dorn-Wallenstein, Neugent, & Levesque 2023; Gangopadhyay et al. 2023; Holmbo et al. 2023; Irani et al. 2023; Jeena, Banerjee, & Heger 2023; Jiang & Liu 2023; Jin, Yoon, & Blinnikov 2023; Kilpatrick et al. 2023b; Könyves-Tóth & Seli 2023; Kuncarayakti et al. 2023; Li, Zhong, & Dai 2023; Lin, Zhang, & Zhang 2023a; Matsuoka et al. 2023; Moore et al. 2023; Moriya et al. 2023c; Park, Yoon, & Blinnikov 2023; Pessi et al. 2023b; Pessi, Prieto, & Dessart 2023; Pursiainen et al. 2023b; Rodríguez, Maoz, & Nakar 2023; Scully, Matzner, & Yalinewich 2023; Shivkumar et al. 2023; Sit et al. 2023; Srinivasaragavan et al. 2023; Stritzinger et al. 2023a, b; Subrayan et al. 2023; Sun, Maund, & Crowther 2023; Teja et al. 2023; Utrobin & Chugai 2023; Van Dyk et al. 2023a; Vasylyev et al. 2023a; Vazquez et al. 2023; Wang et al. 2023a, b; Wen et al. 2023; Williamson et al. 2023; Yan et al. 2023; Zhu et al. 2023; Bjornsson 2023).
I refer here to two explosion mechanisms, the delayed neutrino explosion mechanism and the JJEM. There are many new studies in 2023 of different aspects of the delayed neutrino explosion mechanism (e.g., Aguilera-Dena et al. 2023; Barker, O’Connor, & Couch 2023; Boccioli et al. 2023; Burrows, Vartanyan, & Wang 2023; Gogilashvili, Murphy, & Miller 2023; Navó et al. 2023; Orlando 2023; Powell et al. 2023; Sieverding, Kresse, & Janka 2023; Sumiyoshi et al. 2023; van Baal et al. 2023; Wang & Burrows 2023a, b, 2024; Zha et al. 2023; Maunder et al. 2024; Ghodla & Eldridge 2023). Varma, Müller, & Schneider (2023), for example, find with their 3D simulations that magnetic fields help drive earlier and more energetic explosions (see also Matsumoto, Takiwaki, & Kotake 2024). There is a lower number of new studies of the JJEM (e.g., Antoni & Quataert 2023; Shishkin & Soker 2023; Soker 2023e), including the operation of the JJEM in electron capture supernovae (Wang, Shishkin & Soker, 2024), that potentially can take place for stars with an initial mass of $8.5M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 9.2M_{\odot}$ (e.g., Limongi et al. 2023). I will devote Section 3.2 to describe new results that I consider to strongly support the JJEM, and are toward a breakthrough in answering the question of the CCSN explosion mechanism. I therefore hope this review will motivate more studies of the JJEM.
Both of these explosion mechanisms require instabilities in the zone between the newly born neutron star (NS) and the stalled shock of the collapsing core at $r\simeq 100-150{~{}\rm km}$. The delayed neutrino explosion mechanism requires the instabilities to revive the stalled shock (e.g., Fryer et al. 2023). The JJEM requires the instabilities to increase the angular momentum seed perturbations that the convective zones of the collapsing core supply, in particular the spiral modes of the standing accretion shock instability (SASI). For a recent study of the spiral SASI see Buellet et al. (2023).
Some CCSNe require a central engine to operate after the explosion itself. Many studies attribute the extra energy to a rapidly rotating magnetized NS, i.e., a magnetar (e.g., Acharyya et al. 2023; Bai et al. 2023; Chen et al. 2023c; Gkini et al. 2023; Hu et al. 2023b; Omand & Sarin 2023; Prasanna et al. 2023; Tinyanont et al. 2023; Wang et al. 2023). As I pointed out in the past, the formation process of an energetic magnetar is accompanied by the launching of even more energetic jets, during and/or after the explosion. The majority of magnetar studies continue to ignore the role of jets. The problems of fitting the lightcurves of superluminous CCSNe with magnetars continue with 2023 papers. In Soker (2022) I analyzed the paper by Chen et al. (2023d) who fitted the lightcurves of 70 hydrogen-poor superluminous CCSNe with magnetars, about a third of these have an initial spin period of $P_{\rm mag}\leq 0.002{~{}\rm s}$. However, such rapidly rotating magnetars are likely to be spun up during the last phase of their formation by an accretion disk/belt that launches jets. I argued there that even the slower magnetars in their fitting most likely launched jets at the explosion. In seven out of the 14 CCSNe they fit with ejecta-CSM interaction, the explosion energy is much larger than what the delayed neutrino explosion mechanism can supply, implying a jet-driven explosion. The claim that supernova studies that include energetic magnetars cannot ignore jets continues to hold in 2023. The same arguments hold for the magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that West et al. (2023) and Poidevin et al. (2023) find to fit the hydrogen-poor superluminous SN 2020qlb and SN 2021fpl, respectively, for the number of short-period magnetars that Dong et al. (2023c) argue to power bumpy lightcurves, for those magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that Gomez et al. (2023) and Hinkle, Shappee, & Tucker (2023) fit to hydrogen-poor superluminous CCSNe, and to the magnetar that Fiore et al. (2023) fit to SN 2019neq.
The delayed neutrino explosion mechanism cannot account for energetic CCSNe. Hiramatsu et al. (2023a) find the explosion energy of the type IIn SN 2021qqp to be $E_{\rm exp}\simeq(3-10)\times 10^{51}{~{}\rm erg}$, and Smith et al. (2023) estimate $E_{\rm exp}\simeq(5-10)\times 10^{51}{~{}\rm erg}$ for type IIn SN 2015da. The delayed neutrino explosion mechanism cannot account for such explosion energies, while the JJEM can naturally explain such explosion energies (for a discussion from 2023 see, e.g., Shishkin & Soker 2023). Salmaso et al. (2023) deduced that the explosion energy of SN 2020faa is $\simeq 4\times 10^{51}{~{}\rm erg}$, and discussed the possibility that this peculiar SN II was powered mainly by jets. Likewise, Nagao et al. (2023a) find the explosion energy of the type Icn SN 2021ckj to be $\simeq 4\times 10^{51}{~{}\rm erg}$, an energy the neutrino delayed mechanism cannot account for. Pumo et al. (2023) specifically argue that some of their modeled CCSNe that have high explosion energies challenge the neutrino-driven explosion mechanism. Schweyer et al. (2023) estimate the kinetic energy of SN 2019odp, a type Ib CCSN, to be $E_{\rm exp}\simeq 6\times 10^{51}{~{}\rm erg}$ and note that it is above what the delayed neutrino explosion mechanism can supply. Karamehmetoglu et al. (2023) model several SNe Ibc and find their kinetic energy to be $>3\times 10^{51}{~{}\rm erg}$, more than what the delayed neutrino explosion mechanism can account for. Again, the JJEM easily accounts for such explosion energies. Ergon et al. (2023) discuss the possibility that jets can improve their fitting of the lightcurve of the type IIb SN 2020acat.
Another unsolved problem of the delayed neutrino explosion mechanism is that it seems that it cannot produce the observed mass of ${}^{56}$i in stripped-envelope CCSNe (Sawada & Suwa, 2023). Imasheva, Janka, & Weiss (2023), on the other hand, argue that with proper treatment of the collapsing core, they can produce more ${}^{56}$i, at least for the model appropriate for SN 1987A.
Studies in 2023 continue the exploration of gravitational waves from CCSNe and their detectability (e.g., Bruel et al. 2023; Bugli et al. 2023; Casallas Lagos et al. 2023; Dálya et al. 2023; Heston et al. 2023; Hsieh et al. 2023; Iess et al. 2023; Jakobus et al. 2023; Mezzacappa et al. 2023; Mitra et al. 2023; Pajkos et al. 2023; Pastor-Marcos et al. 2023; Shibagaki et al. 2023; Szczepańczyk et al. 2023; Vartanyan et al. 2023; Villegas et al. 2023; Wolfe et al. 2023; Yuan et al. 2023), including from two cocoons (lobes) that are inflated by fixed-axis energetic jets related to CCSNe (Gottlieb et al., 2023) and by several jet-inflated pairs of lobes (cocoons) in the JJEM (Soker, 2023c). The detection of neutrinos from CCSNe continued to be a topic of interest in 2023 (e.g., Abbasi et al. 2023; Akaho, Nagakura, & Foglizzo 2023; Bendahman et al. 2023; Drago et al. 2023; Fiorillo et al. 2023; Lin et al. 2023b; Mukhopadhyay & Sen 2023; Prasanna, Thompson, & Hirata 2023; Vartanyan & Burrows 2023; Powell et al. 2023). The detection of gravitational waves or neutrinos is not at a stage that allows to support of one explosion mechanism or the other (for the study of CCSN diagnostics see, e.g., Fryer et al. 2023). Therefore, I will not touch on these topics. I will also not touch the subject of axion instability supernova (e.g., Mori et al. 2023).
The presence of CSM around CCSN progenitors and the interaction of the CCSN ejecta with this CSM were major subjects in the 2023 supernova literature (e.g., Akashi, Michaelis, & Soker 2023; Andrews et al. 2023; Ben-Ami et al. 2023; Bietenholz et al. 2023; Bilinski et al. 2023; Bruch et al. 2023; Chugai & Utrobin 2023; Das et al. 2023; Davis et al. 2023; Dessart & Jacobson-Galán 2023; Dickinson et al. 2023; Ercolino et al. 2023; Hiramatsu et al. 2023a; Ishii et al. 2023; Khatami & Kasen 2023; Koo et al. 2023a; Lin et al. 2023; Maeda et al. 2023a, c; Matsuoka & Sawada 2023; Mauerhan et al. 2023; Moran et al. 2023; Moriya 2023; Pearson et al. 2023a; Pellegrino et al. 2023; Pessi et al. 2023a; Petruk et al. 2023; Pursiainen et al. 2023a; Salmaso et al. 2023; Sarmah et al. 2023; Shrestha et al. 2023; Takei et al. 2023; Tsuna, Murase, & Moriya 2023; Tsuna & Takei 2023; Brennan et al. 2023; Murase 2024; Sfaradi et al. 2023), including the role of binary interaction in expelling pre-explosion CSM (e.g., Zenati et al. 2024). [^1] Soker (2013) suggested (based on earlier binary models of similar types of transients) that a sudden pre-CCSN swelling may trigger asymmetric mass loss in binary systems. Five months later Smith & Arnett (2014) repeated this suggestion. Later papers (e.g., Smith 2017; Smith et al. 2023; Bilinski et al. 2023; Mauerhan et al. 2023; Smith et al. 2023) wrongly and consistently have been attributing this suggestion to Smith & Arnett (2014). The ejecta-CSM interaction has become a major topic in relation to the type II SN 2023ixf (e.g., Bersten et al. 2023; Chandra et al. 2023; Guetta et al. 2023; Hiramatsu et al. 2023b; Hosseinzadeh et al. 2023a; Koenig 2023; Liu et al. 2023a; Martinez et al. 2023; Niu et al. 2023; Pledger & Shara 2023; Qin et al. 2023; Sarmah 2023; Sgro et al. 2023; Teja et al. 2023; Van Dyk et al. 2023b; Vasylyev et al. 2023b; Xiang et al. 2023a; Yamanaka, Fujii, & Nagayama 2023; Zimmerman et al. 2023).
Studies found that the ejecta of SN 2023ixf interacted with a CSM that was extended up to a radius of $R_{\rm CSM}\simeq 20-50{~{}\rm AU}$ around SN 2023ixf (e.g., Berger et al. 2023; Bostroem et al. 2023b; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Kilpatrick et al. 2023a; Teja et al. 2023; Smith et al. 2023; Zhang et al. 2023a; Li et al. 2023). In principle, a close CSM might result from an enhanced mass loss rate shortly before the explosion, most likely accompanied by an outburst (eruption; e.g., Tsuna, Takei, & Shigeyama 2023) that might be triggered by activity in the core (e.g. Cohen & Soker 2024). However, several studies find no evidence for pre-explosion outbursts in the years before the explosion of SN 2023ixf (e.g., Dong et al. 2023a; Flinner et al. 2023; Neustadt, Kochanek, & Smith 2023; Jencson et al. 2023; Soraisam et al. 2023; Ransome et al. 2024). This challenges any model where the massive CSM is attributed to impulsive mass ejection years before the explosion of SN 2023ixf. The progenitor of SN 2023ixf did have large amplitude pulsations (e.g., Kilpatrick et al. 2023a; Soraisam et al. 2023). These pre-explosion properties led to the suggestion that instead of a pre-explosion impulsive ejection of a massive CSM, the progenitor of SN 2023ixf maintained an extended zone above its photosphere of rising and falling dense clumps, i.e., the effervescent zone model (Soker, 2023a).
Obergaulinger & Reichert (2023) study nucleosynthesis in fixed-axis jets that power some energetic CCSNe, and find that they have enhanced synthesis of iron-group elements. Leung, Nomoto, & Suzuki (2023) and Leung & Nomoto (2024) study nucleosynthesis in CCSNe driven by fixed-axis jets, including collapsar where the remnant is a black hole. Wanajo (2023) studies nucleosynthesis in neutrino-heated ejecta and neutrino-driven winds, and Friedland, Mukhopadhyay, & Patwardhan (2024) and Psaltis et al. (2023) study the $\nu p$-process in neutrino-driven outflows in CCSNe. These papers do not refer to jets. In general, r-process nucleosynthesis in CCSNe was not a major topic in 2023. Anand et al. (2023), for example, study the lightcurves of broad-line stripped-envelope CCSNe and find no evidence for r-process nucleosynthesis. Nucleosynthesis can be used to study CCSNe. Hoppe et al. (2023) and Liu et al. (2024) present interesting studies of using solar-system isotopes ratio to constrain the stellar properties at the explosion.
Dust formation in CCSN ejecta is another area of intensive research (e.g., Niculescu-Duvaz et al. 2023), as CCSNe might be a significant source of dust in the Universe (e.g., Shahbandeh et al. 2023).
Several papers in 2023 (e.g., Reichert et al. 2023; Volpato et al. 2023) study the rare group of ‘magnetorotational supernovae’, which are driven by fixed-axis jets due to rapid pre-collapse core rotation. While in the delayed neutrino explosion mechanism these rare CCSNe form a separate group (e.g., Fujibayashi et al. 2023a; Maeda, Suzuki, & Izzo 2023), in the JJEM they are just the rapidly rotating tail of CCSNe with slow pre-collapse core rotation (e.g., Soker 2023b). In the fixed-axis case the rapid pre-collapse rotation of the core forces the accretion disk around the newly born NS (and in some cases later around the black hole) to maintain a constant axis, and so are the jets that the accretion disk launches. This leads to a bipolar explosion (e.g., Fang et al. 2023). Vasylyev et al. (2024) measure high continuum linear polarization in the type IIP supernova 2021yja, and discuss the possibilities that it results from CSM interaction or fixed-axis jet explosion. Polarization measurements can teach a lot about the explosion mechanism (e.g., Maund 2023). Nagao et al. (2023b) conclude their spectropolarimetry study of SNe II by stating that “the emergence of a global aspherical structure, e.g., a jet-like structure, might be the key ingredient in the explosion mechanism to produce an energetic SN”. This is compatible with the expectation of the JJEM.
However, such a rapid rotation is rare. On the other extreme, the core rotation is zero, and the angular momentum of the intermittent disks that launch the jets is stochastic such that the jets are fully jittering. The stochastic angular momentum direction results from seed angular momentum perturbations, by the pre-collapse core convection, that are amplified by instabilities above the newly born NS. In the JJEM there is a continuum from CCSNe driven by fully jittering jets, which result from progenitors with very slowly rotating pre-collapse cores, to CCSNe driven by fixed-axis jets that result from rapidly rotating pre-collapse cores. In the frame of the JJEM, this behavior can explain the mass gap between NSs and black holes (Soker, 2023b).
3.2 The breakthrough: evidence for exploding jets
From Section 3.1 it is evident that although the delayed neutrino explosion mechanism is more popular in the literature than the JJEM, it nonetheless encounters some challenges. Here I will argue that the point-symmetric morphologies of three CCSNRs that were studied in 2023 strongly support the JJEM and pose severe problems to the neutrino-driven explosion mechanism.
Late observations, from months post-explosion before the nebular phase (e.g., Haynie & Piro 2023; Medler et al. 2023), through the nebular phase (e.g., DerKacy et al. 2023b; Dessart et al. 2023a, b; Ertini et al. 2023; Fang & Maeda 2023; Liljegren et al. 2023; Margutti et al. 2023; Omand & Jerkstrand 2023; Ravi et al. 2023; Rizzo Smith, Kochanek, & Neustadt 2023; Wesson et al. 2023; Xiang et al. 2023b, and then the remnant phase (e.g., Sato et al. 2023; Zsíros et al. 2023 and more below) can teach us about the progenitor properties and its environment at the explosion time and on the explosion mechanism. Before I describe evidence for jet-driven CCSNe by analyzing the CCSNR phase, I refer to the prediction of the delayed neutrino explosion mechanism that some massive stars collapse into a black hole with no explosion (or with only a very faint explosion). This collapse is termed ‘failed supernovae’.
Some papers in 2023 continue to study this concept of a ‘failed supernova’ (e.g., Coughlin 2023; da Silva Schneider & O’Connor 2023; Disberg & Nelemans 2023; Kuroda & Shibata 2023; Li et al. 2023a; Paradiso et al. 2023). A ‘failed supernova’ is a prediction of the delayed neutrino explosion mechanism where neutrino heating cannot explode the star which in turn collapses to a black hole. On the other hand, in the JJEM there are no failed CCSNe. Even if a black hole is formed there is an explosion, and more likely an energetic explosion (see review Soker 2022). In the JJEM a black hole is likely to be formed when the pre-collapse core is rapidly rotating because this leads to fixed-axis jets that are less efficient in exploding the core (Soker 2023b), i.e., an inefficient jet feedback mechanism (for a recent paper on the accretion disk around the central black hole in CCSNe see Fujibayashi et al. 2023b). Soker (2023b) studies the continuum of explosion outcomes from zero pre-collapse core rotation to rapid pre-collapse rotation in the frame of the JJEM.
The new study by Strotjohann, Ofek, & Gal-Yam (2023) and early results of Byrne & Fraser (2022) question the existence of a large population of ‘failed supernovae’. The detection of a black hole at the center of an SNR, as tentatively claimed by Balakrishnan et al. (2023) for SNR G351.9-0.9, shows that black holes can form in CCSNe, rather than in ‘failed supernovae’. The detection by Beasor et al. (2023) of a luminous infrared source at the position of the ‘failed supernova’ candidate N6946-BH1 shows, as they claim, that it better fits a type II intermediate luminosity optical transient (type II ILOT) than a ‘failed supernova’ (but see Kochanek 2023; Kochanek, Neustadt, & Stanek 2023). The above-cited studies challenge the prediction of the delayed neutrino explosion mechanism of ‘failed supernovae’. I reiterate my claim that studies of CCSNe that form black holes cannot ignore jets.
Let me then turn to a prediction of the JJEM that in some (but not all) cases the jittering jets imprint a point-symmetric morphology. Most of the jittering jets during the explosion process, typically lasting for about one to several seconds and more, deposit their entire energy in the core of the star and explode it. From that point on the explosion is similar (but not identical) to that of the delayed-neutrino explosion mechanism. However, in some cases, the few last pairs of jets that are launched after the inner parts of the stars are exploded might leave an imprint on the outer ejecta. In most cases only the last pair leaves an imprint, forming two opposite ‘ears’ (e.g., Soker 2023d). An ear is a protrusion from the main ejecta with a decreasing cross-section with distance from the center. In a minority of cases, more than one pair of jets might leave imprints, forming a point-symmetric morphology.
In 2023 I identified point-symmetric morphologies in three CCSNRs. In Section 1 I discussed the point-symmetric morphology and why jets are most likely to shape such morphologies.
In Soker (2023f) I studied the morphology of the CCSNR N63A based on new X-ray images that Karagöz et al. (2023) present. I present the image and my identification of a point-symmetric morphology in Figure 1. I identified the point-symmetric morphology to be the three pairs of ears that are marked in the figure. This point symmetry cannot be shaped by ejecta-CSM interaction or ejecta-ISM interaction (Section 1). Therefore, this point-symmetry strongly supports the JJEM. Moreover, the differences in sizes of the two ears in each pair are accounted for by the JJEM.
Figure 1: The CCSNR N63A has a point-symmetric morphology. A Chandra X-ray image of SNR N63A (logarithmic-scale $0.3-7.0{~{}\rm keV}$). The green lines and the small-yellow writings are from the original figure of Karagöz et al. (2023). The red-solid lines, the dashed-blue lines, and the red and blue writing are additions from Soker (2023f) that identify a point-symmetric morphology.
In each pair, one ear is significantly larger than the opposite ear. This is not due to an asymmetrical medium on sides A and B (marked in the figure) as Ear 1a is larger than Ear 1b, but Ears 2a and 3a are smaller than Ears 2b and 3b, respectively. I, therefore, concluded (Soker, 2023f) that the asymmetry of each pair results from asymmetrical opposite jets at the time of the jets’ launching. This asymmetry likely results from that the accretion disk that launches the jets in the JJEM has no time to fully relax. In the JJEM there are $\approx{\rm few}$ to $\approx 30$ jet-launching episodes, each lasting for $\approx 0.01-0.1{~{}\rm s}$. The viscous timescale of the accretion disk in each jet-launching episode is not much shorter than this timescale. Therefore, the disk has no time to fully relax to a thin accretion disk. Since the instabilities that must operate in the JJEM are likely to form each accretion disk with unequal sides of the temporarily equatorial plane, this asymmetry lasts for most of the jet-launching episode. The two unequal sides of the accretion disk are very likely to launch two opposite jets that have different energy, mass, and velocity. These asymmetrical two opposite jets in a pair will inflate unequal ears.
In Soker (2023d) I classified 14 CCSNRs into five groups according to their morphology as shaped by jets. I argued that the degree of pre-collapse core rotation is one of the main parameters that determines the morphological class. The faster the pre-collapse core rotation is the less random the directions of the jets in the different jet-launching episodes are. For a very slow (or zero) pre-collapse core rotation, the jets jitter almost randomly. This, I suggested, might lead to a point-symmetric morphology with pairs of opposite structures in many directions. In that study, I identified a point-symmetric morphology in the Vela SNR. The Vela SNR is well-studied (e.g., Mayer et al. 2023; Ritchey 2023; Kargaltsev, Hare, & Lange 2024). I present the identification of the point-symmetric morphology in Figure 2 taken from Soker (2023d). Earlier studies already noted the existence of pairs of opposite structural features, as marked in the figure. However, only in Soker (2023d) did I explore the full point symmetry and connect it to jittering jets that exploded the progenitor of the Vela SNR.
Figure 2: The point-symmetric morphology of the Vela SNR, as identified by Soker (2023d). The ROSAT X-ray image is from Aschenbach et al. (1995), and the figure is adapted from figure 1 of Sapienza et al. (2021). Clumps A to F were identified by Aschenbach et al. 1995. The white AG-line and the labeling of the clumps are by Sapienza et al. (2021). The thick-yellow FJ-line and DE-line and the two dashed-black lines are marked by Soker (2023d). Each of the two dashed-black lines connects a clump to an assumed counter jet.
SN 1987A has been one of the prime focuses of the SN community since its explosion (for 2023 papers see, e.g., Arendt et al. 2023; Cikota et al. 2023; Dohi et al. 2023; Fiorillo et al. 2023; Jones et al. 2023; Kangas et al. 2023; Larsson et al. 2023; Weishi Li et al. 2023; Sapienza et al. 2023). I am interested here in the ejecta morphology and its implication on the explosion mechanism. Ono et al. (2023) consider a bipolar explosion morphology for SN 1987A. They do not refer to jets though.
What allows me to (tentatively) claim for the JJEM for 1987A is the new observation of SN 1987A by JWST (Arendt et al. 2023) as I present in Figure 3. All marks on the original figure are from Soker (2024b), where more details can be found. In that paper, I identified a tentative point-symmetric morphology in the new JWST image of SN1987A. As said, the JJEM predicts that in some CCSNe the last jets imprint a point-symmetrical structure. I suggest such a structure in the ejecta of SN 1987A, and suggest that it was exploded by jittering jets. The asymmetry within each pair of two clumps is again attributed to the not relaxed intermittent accretion disks that launched the jets during the explosion.
Figure 3: An image of SN 1987A from JWST [source: NASA Webb Telescope Team, AUG 31, 2023; NASA, ESA, CSA: Matsuura, M. (Cardiff University), Arendt, R. (NASA’s Goddard Spaceflight Center & University of Maryland, Baltimore County), Fransson, C.]. [^3] For the original image see: https://www.nasa.gov/missions/webb/webb-reveals-new-structures-within-iconic-supernova/ Red represents emission at 4.44 microns, orange at 4.05 microns, yellow at 3.23 microns, cyan at 1.64 and 2.0 microns, and blue at 1.5 microns. The arrows in the left panel are additions from Soker (2024b). These arrows point at clumps, namely, some small areas in the outer regions of the blue-cyan structure that are brighter than their surroundings. The three black lines in the right panel, also from Soker (2024b), connect pairs of clumps that the arrows in the left panel point at. The red double-headed arrow indicates the direction of the elongated axis of the blue-cyan structure. The two red-dashed lines signify a possible structure of two opposite jets.
Let me end with the study of the northeast jet of the SNR Cassiopeia A by Koo et al. (2023b). They conclude that both the neutrino-driven explosion mechanism and the magnetocentrifugal jet-induced explosion (fixed-axis jets) have problems in accounting for the properties of the northeast jet of Cassiopeia A. This is a known result. As I commented in the past, the JJEM can best explain the properties of Cassiopeia A (e.g., Soker 2017).
I consider the point-symmetric morphologies of these three CCSNRs (and a speculative possible identification of a point-symmetric morphology in SNR CTB 1; Bear & Soker 2023), supported by earlier studies of CCSNR morphologies, to be a step toward a breakthrough in the determination of the CCSN explosion mechanism. The JJEM cannot be ignored in analyzing CCSNe.
4 Type Ia supernova ejecta
4.1 SN Ia scenarios
The diversity of normal and peculiar SNe Ia (e.g., DerKacy et al. 2023c; Desai et al. 2023b; Dwarkadas 2023b; Fausnaugh et al. 2023; Graur et al. 2023; Harris et al. 2023; Harvey et al. 2023; Hinse et al. 2023; Holland-Ashford et al. 2023; Huang et al. 2023; Keegans et al. 2023; Larison et al. 2023; Li et al. 2023b; Lu et al. 2023; Nascimento, França, & Reis 2023; O’Brien et al. 2023; Ogawa, Maeda, & Kawabata 2023; Pearson et al. 2023b; Peterson et al. 2023; Sharon & Kushnir 2023; Singh et al. 2023; Tiwari et al. 2023; Toy et al. 2023; Xi et al. 2023; Yang et al. 2023a; Graur 2024) have motivated the consideration of all major SN Ia scenarios in 2023. I classify six scenarios as in Soker (2024a), and present them in Table 1 (for a different partition of the scenarios, in particular sub-classification of the SD scenario, see the table in the review by Liu, Röpke, & Han 2023).
Table 1: SN Ia scenarios and their ability to form point-symmetric or spherical SNRs
Group $N_{\rm exp}=1$: Lonely WD $N_{\rm exp}=2$
Outcome Mostly normal SNe Ia Mostly peculiar SNe Ia
SN Ia Scenario
Core Degenerate
Double Degenerate - MED
Double Degenerate
Double Detonation
Single Degenerate
WD-WD collision
Name
CD
DD-MED
DD
DDet
SD-MED or SD
WWC
MED time
MED
MED
0
0
MED or 0
0
$\mathbf{[N_{\rm sur},M,Ej]}$${}^{[{\rm 2}]}$
$[0,M_{\rm Ch},{\rm S}]$
$[0,M_{\rm Ch},{\rm S}]$
[…]
$[1,M_{\rm Ch},{\rm S~{}or~{}N}]$
[…]
Point symmetry in the SNR
Expected in SNIPs with point-symmetric PNe. SN 1181; G1.9+0.3
Very rare: A SNIP with point-symmetric PN. SN 1181
Very rare: A SNIP with point-symmetric PN.
Extremely rare.
Possible: a symbiotic progenitor; Low-mass CSM.
Extremely rare; large-scale departures from ellipticity.
Notes: (1) Cells in bold red indicate new suggestions/speculations from 2023 that I support in this review. The two SNRs Ia that I focus on are the peculiar SN 1181, which I suggest might result from the DD-MED or the CD scenarios, and SNR G1.9+0.3, which results from the CD scenario in a planetary nebula (SNIP). (2) Abbreviation. MED time: Merger to explosion delay time (includes mass transfer to explosion delay time). $N_{\rm exp}$ is the number of stars in the system at the time of the explosion. $N_{\rm sur}=1$ if a companion survives the explosion while $N_{\rm sur}=0$ if no companion survives the explosion; in some peculiar SNe Ia, the exploding WD is not destroyed and it also leaves a remnant, i.e., $N_{\rm sur}=2$. $M_{\rm Ch}$ for near-Chandrasekhar-mass explosion; sub-$M_{\rm Ch}$ for sub-Chandrasekhar mass explosion. The ejecta morphology Ej: S for scenarios that can lead to a spherical SNR while N indicates that the scenario cannot form a spherical SNR.
[…]
In the double degenerate (DD) scenario (e.g., Dimitriadis et al. 2023; Karthik Yadavalli et al. 2023; Moran-Fraile et al. 2023; Qi, Liu, & Wang 2023; Srivastav et al. 2023a) and in the DD-MED scenario, two WDs merge after losing angular momentum to gravitational waves. In the DD scenario, the explosion occurs within a few dynamical timescales, like in the violent merger channel, and before the non-spherical imprints of the binary system disappear. This DD scenario channel has no MED time and the explosion is non-spherical. Practically there are two stars at the explosion, i.e., $N_{\rm exp}=2$. In the DD-MED scenario channel, the explosion takes place only after the merger remnant relaxes. It is likely to be a spherical near-$M_{\rm Ch}$ explosion. At explosion, there is only one star (as wide companions play no roles), i.e., $N_{\rm exp}=1$.
The double detonation (DDet) DDet scenario, where a helium layer accreted from a companion is ignited and detonates the WD, was relatively popular in 2023 (e.g., Collins et al. 2023; Zingale et al. 2023) with attribution of this scenario to some specific peculiar SNe Ia (e.g., Liu et al. 2023b, c; Padilla Gonzalez et al. 2023a, b for four peculiar SNe Ia, respectively). Several studies attribute hyper-runaway/hypervelocity WDs to the WD helium-donor companion in the DDet scenario (e.g. El-Badry et al. 2023; Werner et al. 2023). However, new studies question whether this scenario can explain normal SNe Ia. Igoshev, Perets, & Hallakoun (2023) argue that the numbers and properties of hypervelocity WDs with velocities of $>1000{~{}\rm km}{~{}\rm s}^{-1}$ might at most account for a small fraction of SNe Ia, but might account for some peculiar SNe Ia. Braudo & Soker (2023) study the dynamics of the surviving WD and of the ejecta, and conclude that very fast runaway WDs require massive exploding CO WD. This makes such a scenario rare and supports the claim that the DDet scenario leads mostly to peculiar SNe Ia but not to normal SNe Ia.
In the single degenerate (SD) scenario (e.g., Hosseinzadeh et al. 2023b; Li, Liu, & Wang 2023, and Ruiz-Lapuente 2023 and Ruiz-Lapuente et al. 2023 for Tycho SNR) a WD accretes mass from a non-degenerate companion and explodes when reaching a near-Chandrasekhar mass. It can explode as it reaches the exploding mass, or there might be a long delay time from mass accretion to explosion (MED time). The non-detection of surviving companions in SNRs Ia, e.g., Shields et al. (2023) for SNR 0509-67.5, Tucker & Shappee (2023) for SN 2011fe, and Rodriguez, Yanza, & Dzib (2023) for Tycho’s SNR, continues to be a major challenge to the SD scenario (and also to the DDet scenario). Considering the difficulties of the SD scenario, I find it unjustified that in 2023 many papers mention only the SD and the DD scenarios, ignoring other SN Ia scenarios.
The WD-WD collision (WWC) scenario, where the collision of two WDs detonates both of them (e.g., Glanz, Perets, & Pakmor 2023), has been shown over the years to be able to account for no more than a minor fraction of all SNe Ia (see earlier reviews cited in Section 1).
Triple star interaction might enhance the rate of each of these scenarios (e.g., Rajamuthukumar et al. 2023). Some studies do not support any specific scenario over another (e.g., Meng et al. 2023).
Some channels are aimed at peculiar SNe Ia. Zenati et al. (2023) construct a channel of the DD scenario where the disruption of a low-mass carbon–oxygen (CO) WD by a hybrid helium–CO (HeCO) WD forms calcium-rich peculiar SN Ia and leaves a WD remnant. Srivastav et al. (2023b) proposed the DD scenario for SN 2022ilv, a super-$M_{\rm Ch}$ peculiar SN Ia.
Studies in 2023 have continued to explore both sub-Chandrasekhar (sub-$M_{\rm Ch}$) and near-Chandrasekhar ($M_{\rm Ch}$) explosions (e.g., Chakraborty et al. 2023; DerKacy et al. 2023a, b; Guo et al. 2023; Hoeflich et al. 2023; Leung & Nomoto 2023; Maeda et al. 2023b; Ni et al. 2023a; Yarbrough et al. 2023).
The existence of CSM around some SNe Ia continues to be puzzling (e.g. Dwarkadas 2023a; Mao-kai, Li-fan, & Xiao-feng 2023; Moriya et al. 2023b). However, it is not questionable anymore that some SNe Ia have massive CSM, as Guest et al. (2023) find for the SNR 0519–69.0 in the Large Magellanic cloud. Such is SNR G1.9+0.3 that I review in more detail in Section 4.2.3.
The excess emission at early SN Ia phases (within a few days) is still puzzling (e.g., Lim et al. 2023; Ni et al. 2023a, b; Wang et al. 2023c). Processes might include the collision of the ejecta with a companion, with a close CSM, or the presence of ${}^{56}$i in the outer zones of the ejecta. Each of these has implications for the possible SN Ia scenarios. Wang et al. (2023c) conclude that no current model can adequately explain the full set of observations of the early emission excess of SN 2023bee. It seems that the early light excess cannot yet point to one leading SN Ia scenario, but has the potential to do so with more observations and their analysis.
The observation of SNe Ia during the nebular phase (e.g., Blondin et al. 2023; Camacho-Neves et al. 2023; Chen et al. 2023a; Deckers et al. 2023; Kumar et al. 2023; Kwok et al. 2023a; Liu et al. 2023d; Ni et al. 2023a) is critical to reveal the explosion mechanism as it can, for example, reveal the total mass of some isotopes that are nucleosynthesis at the explosion and their velocity distribution. Chen et al. (2023a) and Kwok et al. (2023a) deduce that SN 2021aefx has a spherically symmetric distribution of observed isotopes at the nebula phase. This supports the notion that many normal SNe Ia have large-scale spherical explosions (for review see Soker 2019). Blondin et al. 2023 argue that no existing SN Ia explosion model can fit all observed properties of SM 2021aefx.
The relation between SN Ia types and their environment is another promising direction to learn about SN Ia scenarios, normal and peculiar (e.g., Nugent, Polin, & Nugent 2023). Some peculiar SNe Ia might come from a different population than normal SNe Ia. Barkhudaryan (2023) for example, studies the height of SNe Ia and argues that 91T-like peculiar SNe Ia (which are at the slow-declining, hot, luminous end of normal SNe Ia) originate from relatively younger progenitors with ages of $\approx 100{~{}\rm Myr}$, 91bg-like peculiar SNe Ia (which are fast-declining, cool, subluminous peculiar SNe Ia) arise from significantly older progenitors of ages of $\approx~{}10{~{}\rm Gyr}$, and normal SNe Ia are from progenitors of $\approx 1-10{~{}\rm Gyr}$.
In analyzing observations one must bear in mind all SN Ia scenarios. This is evident in the study of SN 2020eyj. Kool et al. (2023) observed SN 2020eyj that they found to be an SN Ia that interacts with a helium-rich CSM. They attribute SN 2020eyj to the SD scenario with a helium non-degenerate donor star. Kool et al. (2023) specifically write that their observation “leave the SD scenario as the only viable alternative”. However, Soker & Bear (2023) propose a new channel of the CD scenario where there are two major CEE phases. In the second CEE phase, the helium-rich core expands to interact with a WD companion, forming a helium-rich CSM before the explosion. The explosion takes place with a MED time of weeks to tens of years.
4.2 The group of lonely white dwarf scenarios
One of my claims in this review is that many (but not all; see below) of the peculiar SNe Ia result from channels that have two stars at the explosion $N_{\rm exp}=2$, as I present in Table 1, while normal SNe Ia result mainly from the lonely WD group (e.g., Braudo & Soker 2023), namely, $N_{\rm exp}=1$. I concentrate here on two SNRs that were analyzed in 2023, SN 1181 which is a peculiar SN Iax, and SNR G1.9+0.3, which is probably a normal SN Ia. I first discuss some non-lonely WD scenarios.
4.2.1 Peculiar SNe Ia from $N_{\rm exp}=2$ scenarios
I present some recent studies that further motivate me to suggest that most peculiar SNe Ia result from $N_{\rm exp}=2$ scenarios, i.e., scenarios that have two stars at the time of the explosion.
From their observation of the super-Chandrasekhar peculiar SN Ia (03fg-like) SN 2020hvf during the nebular phase Siebert et al. (2023a) conclude that the explosion is asymmetrical with two velocity components. They discuss the DDet scenario in a double-degenerate system. Namely, one WD exploded and one survived. This suggestion, if holds, shows that a scenario with two stars at explosion can lead to an aspherical peculiar SN Ia. That paper also shows the potential of observations during the nebular phase in revealing the exploding system.
Kwok et al. (2023b) demonstrate the importance of observations during the nebular phase in their study of the peculiar (03fg-like or super-Chandrasekhar) SN 2022pul. The presence of carbon-oxygen rich CSM (Siebert et al., 2023b), the composition of the ejecta, and the asymmetry of the ejecta led Kwok et al. (2023b) to conclude that the best model of this peculiar SN Ia is the violent merger channel of the DD scenario. This channel has two WDs at the explosion, $N_{\rm exp}=2$. Li et al. (2023c) suggest the violent merger channel of the DD scenario for the peculiar SN 2016ije. In a different study of ‘super-Chandrasekhar’ SNe Ia, Fitz Axen & Nugent (2023) find the DD scenario to better fit these peculiar SNe Ia than an explosion of a massive magnetized WD.
The study by Hoogendam et al. (2023) is very relevant to the proposition that peculiar SNe Ia have different progenitors than most normal SNe Ia. Hoogendam et al. (2023) find that 2002es-like and 2003fg-like peculiar SNe Ia have different pre-peak UV color types of evolution compared to normal SNe Ia. They further suggest that these peculiar SNe Ia have different progenitors, and propose that these peculiar SNe Ia may originate in low-metallicity DD or CD scenarios enshrouded by a carbon-rich circumstellar medium.
4.2.2 SN 1181 and its SNR Pa 30
Several papers in 2023 revealed the nature of Pa 30 as the SNR of the historical SN 1181, which was an SN Iax (Fesen, Schaefer, & Patchick, 2023; Ko et al., 2023; Lykou et al., 2023; Ritter et al., 2023). The large-scale circular morphology of Pa 30 with a hot remnant at its center is critical to the understanding of SN 1181. Fesen, Schaefer, & Patchick (2023) conduct a thorough study of Pa 30. They note that the highly spherical morphology of SNR Pa 30 is in sharp contrast to the strongly asymmetric SN Iax ejecta in some models. They attribute the filamentary structure of the round SNR to the photoionization effects of clumpy ejecta (as the modeling of some filamentary planetary nebulae). I refer here only to the large-scale spherical morphology. I, therefore, present one image from Fesen, Schaefer, & Patchick (2023) in Figure 4.
Figure 4: An [S II] image of SNR Pa 30 from Fesen, Schaefer, & Patchick (2023). This image presents the large-scale spherical morphology of SNR Pa 30, the remnant of SN 1181. It is an SN Iax with a hot remnant at its center.
Schaefer (2023) presents a detailed discussion of SN 1181 and its SNR Pa 30. He proposes that the best scenario to match observations is a merger of oxygen-neon (ONe) and carbon-oxygen (CO) WDs (see also Ko et al. 2023), like in the simulations by Kashyap et al. (2018). However, the simulations by Kashyap et al. (2018) show that the outcome is a highly non-spherical ejecta, contrary to the structure of Pa 30. The requirement from scenarios of normal SNe Ia to be able to form spherical SNRs was the highlight of an earlier review of SNe Ia (Soker, 2019). This requires the system to relax after a binary interaction, whether a full merger or a mass transfer process. Namely, there must be a merger to explosion delay (MED) time. Maguire et al. (2023) study SN 2020udy which is another SN Iax and argue that it was an incomplete explosion of a near-$M_{\rm Ch}$ WD. For SN 2020udy, I would also claim for a MED time. The importance of MED time in normal SN Ia scenarios is the highlight of another earlier review (Soker, 2018).
Bogomazov & Tutukov (2023) study the DD scenario with CO WD and one ONe WD for normal SNe Ia (for different outcomes of the merger of a CO WD with a ONe WD, including a possible electron capture CCSN, see Wu et al. 2023). However, SN 1181 is a peculiar SN Ia. This and its spherical ejecta suggest a lonely WD at the explosion time, which resulted from either the CD scenario where either the core of the asymptotic giant branch star or the WD companion, which merged with the core, being ONe rich (a scenario that was studied by Canals, Torres, & Soker 2018). Another possibility is the DD-MED scenario where a CO WD and an ONe WD merged to form a near-$M_{\rm Ch}$ WD that exploded after it relaxed, hence the spherical explosion.
I conclude that SN 1181 with its large-scale spherical SNR suggests that it is a result of one of the lonely WD scenarios (Table 1).
4.2.3 SNR G1.9+0.3
In Soker (2024a) I identified a point-symmetry in the morphology of SNR G1.9+0.3, the youngest known SNR in the Galaxy, which exploded around 1890-1900. I analyzed the X-ray image of this SNR that Enokiya et al. (2023) released and pointed at ten pairs of opposite clumps or filaments in the X-ray morphology, in addition to the opposite pair of ears that was noticed in the past. I present the analysis in Figure 5 with more details in the caption. Known SN Ia explosion models do not form such point-symmetric structures. I therefore concluded that the point-symmetric structure comes from the CSM which was a planetary nebula. Namely, SNR G1.9+0.3 is an SNIP, i.e., an SN inside a planetary nebula. In Soker (2024c) I discussed SNR G1.9+0.3 together with the brightest planetary nebulae in the Galaxy and proposed that the planetary nebula into which the progenitor of SNR G1.9+0.3 exploded had a multi-polar morphology.
Figure 5: The point-symmetric morphology of type Ia SNR G1.9+0.3. The original X-ray image with contours of CO emission and the notations of the ellipse and coordinates are from Enokiya et al. (2023). The double-headed arrows (DHA; red, white, and yellow), the notation of the ears, and the inset on the right panels with the notations of rotation and bending are from Soker (2024a). The left panel presents the basic point-symmetric morphology. Each of the lines DHA-a to DHA-f points at a pair of clumps (twin clumps). The group of twin pairs and the pair of ears form the point-symmetric morphology. Line DHA-$\tau$ is a tentative pair because the clumps are small and faint. The pairs of clumps that two white double-headed arrows point at are treated differently on the right panel. The right panel focuses on the two opposite ears and the arc at the base of each ear. The closed dashed-white line on the upper arc marks the boundary of the arc and the dashed black lines mark the three emission peaks within the arc. The boundary of the arc with its three peaks is rotated around itself by $180^{\circ}$ and matched to the arc at the base of the bottom (eastern) ear by a $9^{\circ}$ displacement as the inset indicates. Namely, there is a $9^{\circ}$ bend point symmetry of the two ears (DHA-0) and of the arc at the base of each ear (DHA-1 to DHA-3).
[…]
I also note that Kobashi et al. (2023) conclude that the CSM mass of Tycho’s SNR is larger than expected in the SD scenario and better matches the CD scenario. Tycho large-scale circular morphology also suggests a scenario from the lonely WD group of scenarios.
The conclusion of the study of SNR G1.9+0.3 is that it was an explosion inside a planetary nebula of a near-$M_{\rm Ch}$ lonely WD, most likely in the frame of the CD scenario, with a tiny probability for the DD scenario.
4.2.4 Lonely WDs at the explosion time
For its novelty, the group of lonely WD scenarios deserves this subsection. As I indicate in Table 1 the group of lonely WD scenarios includes the DD-MED scenario and the CD scenario. The exact definition of a lonely WD explosion is one where at the time of explosion there is only one star in the system and it is dynamically relaxed. This implies that the explosion might be spherical, or axi-symmetrical if the lonely WD is rapidly rotating. No other material explodes but the WD itself, nor a companion nor an accretion disk.
Observed properties are more complicated and not all explosions are spherical. For example, Ni et al. (2023a) deduce that SN Ia 2021aefx is an $M_{\rm Ch}$ explosion but the explosion is not spherical. In the lonely-WD scenarios, this is accounted for by an explosion of a rapidly rotating WD or by an off-center deflagration to detonation transition. The rotating WD is likely to sustain a strong magnetic field that might account for the high polarization of some emission lines of the descendant SN Ia. Yang et al. (2022) observed high polarization in the Ca II near-infrared triplet of the normal SN Ia 2021rhu, and propose that magnetic field is the cause of this high polarization.
Consider first the DD-MED scenario. The dynamical relaxation timescale of a WD remnant of a merger is about several seconds. An accretion disk around the WD should also be depleted, which can take several minutes. Therefore, the MED time should be several minutes or longer. The accretion disk can blow a wind or launch jets forming a disk-originated matter (DOM). If the MED time is shorter than about a few days, the ejecta collides with the DOM within hours to a few days, leading to an early excess emission (e.g., Srivastav et al. 2023a for SN 2022ywc). Although the DOM might have a bipolar structure, the explosion can be spherical. If the MED time in the DD-MED scenario is much longer, then we can have a spherical explosion of a normal SN Ia, or a spherical mass ejection in a peculiar SN Ia that leaves a WD remnant after the explosion. The latter channel is a possible scenario for SN 1181 that formed the spherical SNR Pa 30 (Section 4.2.2).
In the CD scenario, the MED time is longer than several days because the hydrogen-rich envelope that was ejected during the common envelope evolution must be at large distances for the explosion to be classified as SN Ia. In rare cases, the last envelope to be ejected is helium-rich (Soker & Bear, 2023). The CD scenario can explain SNe Ia that have massive CSM, up to $M_{\rm CSM}\approx 3M_{\odot}$. Since this CSM is or was ionized by the WD merger remnant, it is or was a planetary nebula at the time of the explosion, i.e., the SN Ia is an SNIP. As with many other planetary nebulae that are formed by binary interaction, the CSM might be elliptical or bipolar and possess a point-symmetric morphology. I suggest an SNIP in the CD scenario for SNR G1.9+0.3 (Section 4.2.3). The CD scenario with a MED time of $>10^{6}{~{}\rm yr}$, such that the CSM does not exist anymore, might also account for SN 1181 and its Pa 30 remnant.
5 Summary
5.1 Point symmetric supernova remnants (SNRs)
I used the morphologies of CCSNRs and SNRs Ia as a common ground to include both CCSNe and SNe Ia in the same review of SN papers in 2023.
The research of SNRs is a major topic with tens of papers in 2023 (Section 2) that cover many aspects of SNRs. I concentrated mainly on SNR morphologies, and in particular on point-symmetric morphologies. As I described in Section 2, the point-symmetric morphology requires that either the explosion itself, as I argue for CCSNe, and/or the CSM, as I argue for SNe Ia, have point-symmetric morphology to begin with. The interaction of a non-point symmetric explosion with a non-point-symmetric CSM or ISM, cannot form point-symmetrical SNRs.
Another common ground for CCSNe and SNRs is that observations during the nebular phase, when the entire ejecta can be observed, might reveal the explosion mechanism and progenitor properties.
5.2 Core collapse supernovae (CCSNe) in 2023
Tens of papers in 2023 further demonstrate the diverse properties of CCSNe. Some major topics that research papers of CCSNe in 2023 cover are (Section 3.1) the presence of CSM around CCSNe and the ejecta-CSM interaction, the role of magnetars in powering superluminous CCSNe, the effects of rapidly rotating pre-collapse cores, and the process of nucleosynthesis in CCSNe. SN 2023ixf is a major CCSN of 2023. Many CCSNe have dense CSM at the time of the explosion that is ejected in an outburst shortly before the explosion. The study of SN 2023ixf that had no pre-explosion outburst forces the consideration of a pre-collapse long-lived extended dense zone above the photosphere, like the effervescent zone model (Section 3.1), as an alternative in some cases.
Many papers study the detection of gravitational waves and neutrinos from CCSNe, although these are not yet in a stage to constrain explosion mechanisms.
In Section 3.2 I described the analysis of three CCSNRs from 2023 and further discussed the delayed neutrino explosion mechanism and the JJEM.
I summarize the main claims of this review in regard to the explosion mechanism of CCSNe as follows.
  1. 1.
    The inclusion of energetic magnetars in CCSNe must be accompanied by a discussion of the role of jets because jets accompany the formation of energetic magnetars (Section 3.1).
  2. 2.
    The delayed neutrino explosion mechanism predicts ‘failed supernovae’, which are not supported by observations (Section 3.2). Studies of CCSNe that form black holes cannot ignore jets that, according to the JJEM, explode the star even when the remnant is a black hole.
  3. 3.
    According to the JJEM some CCSNRs might have point-symmetric morphologies. This was verified in three CCSNRs in 2023 (Figures 1 - 3). I consider this group of three-point-symmetric CCSNRs (and a speculative identification of a point-symmetric morphology in SNR CTB 1; Bear & Soker 2023) to constitute a breakthrough in our quest for the explosion mechanism of CCSNe, which, I suggest, is the JJEM.
5.3 Type Ia supernovae (SNe Ia) in 2023
In Section 4 I reviewed the scenarios of normal and peculiar SNe Ia that have been studied in 2023. I presented the classification into six scenarios in Table 1 and criticized studies that even in 2023 still mention only the SD and DD scenarios and ignore other scenarios, in particular the CD and DD-MED scenarios.
The main conclusions related to SNe Ia are as follows.
  1. 1.
    The diversity of both normal and peculiar SNe Ia suggests that more than just two scenarios are involved in forming the majority of all types of normal and peculiar SNe Ia (likely 4 to 5 scenarios). Studies should be open to the different scenarios (Table 1). The habit of mentioning only the SD and DD scenarios should be abandoned.
  2. 2.
    Studies in 2023 have shown that observations during the nebular phase could deduce that some peculiar SNe Ia result from some specific scenarios, mainly, but not only, from non-lonely scenarios, i.e., $N_{\rm exp}=2$ (Sections 4.1 and 4.2.1).
  3. 3.
    As noticed by past reviews of SNe Ia, the large-scale spherical morphologies of many SN Ia remnants constrain the common SN Ia scenario(s) to be able to avoid aspherical effects, such as a close companion. The large-scale spherical structure of Pa 30, the remnant of the peculiar SN Ia 1181 (Section 4.2.2), and other considerations in 2023 have motivated the introduction of the group of lonely WD scenarios (Table 1).
  4. 4.
    The point symmetric morphology of SNR G1.9+0.3 is the strongest ever argument for SN Ia inside a planetary nebula (SNIP; Section 4.2.3). Together with SNR Pa 30 of SN 1181, they also put the SN Ia lonely WD scenarios, which have MED time, on very solid ground.
  5. 5.
    Taking the above conclusions, the group of lonely WD scenarios (and non-lonely WD scenarios) should be a major lead in our quest to related SN Ia scenarios (Table 1) with observed normal and peculiar SNe Ia.
More research, like the search and detection of supernova progenitors, both CCSNe (e.g., Strotjohann et al. 2023; Van Dyk et al. 2023) and SNe Ia, are essential to verify or reject the breakthroughs I pointed out in this review. Essential to the study of explosion mechanisms and progenitor properties is to consider both theoretical explosion mechanisms of CCSNe and all SN Ia scenarios.
Acknowledgements
I thank Hila Glanz and Amit Kashi for the useful comments. In writing this review I was motivated by the results and discussions of two 2023 meetings on supernovae, SuperNova EXplosions Conference (SNEX, Technion, Israel; chaired by Hagai Perets and Yossef Zenati) and SuperVirtual 2023 (virtual; chaired by Melina Cecilia Bersten and Takashi Moriya). This research was supported by grants from the Pazy Research Foundation and a grant from the Israel Science Foundation (769/20).